Project

General

Profile

Actions

Meeting 2014-12-17

Date: 17th December 2014
Time: 15:00 UK time
Attendance: Craig, Geoff, Ed, Ray, David, Adam, Pierrick
Apologies:

Agenda and Minutes

0. Actions from Last Meeting

  • Check with R. Bayes that this does not affect things in main CDB geometry - C. Heidt - Polished up geometry from C Hunt need to get files from Geoff. - New action GB and MU - Geoff and Melissa meet with Jason tomorrow.
  • Ensure central fibre is correctly defined in geometry and with Tracker group and correct if necessary - C Heidt CARRIES OVER
  • Issues found with NIM paper description of external waveguide end to end mapping check with Geoff and SW - MU AT MEETING tomorrow with Geoff
  • installation manual plan - GB MU Ongoing

1. Software: AD, ES, CH

  • Schedule - DA

1.1 Documentation - DA, AD, ES, C. Heidt

1.2 Geometry, Calibrations, Configurations, CDB - C. Heidt

1.3 Monte Carlo

  • MC Noise - C. Heidt
  • ADC smearing needs fixing - C. Heidt
*Progress made on noise and detector response, but needs to model individual channels and light yield. This will accurately define inefficiencies.
  • Chris Hunt is taking over Ed Santos's Kalman work
  • Kalman doesnt work on straights. Chris and Adam looking to fix

1.4 Pattern Recognition - AD, DA

  • Low pt

1.5 Online Displays - MU Ongoing

1.6 Performance and Analysis

  • Emittance error - discrepancy by C. Hunt results and ES results (ES ~0.6%, CH ~ 0.8 - 1.5%)
  • Bias difference between trackers to be investigate - C. Hunt, ES
  • Spacepoint efficiency study - C. Heidt, AD
  • Low pt performance - AD
  • Efficiency and Purity figures for different beams and geometries

1.7 Issues

1.8 Tracker Software Paper - AD

Published as MICE note and referred to in MPB report

2. DAQ/Calibration/Controls - DA, EO, CmW

  • Tracker DAQ / electronics QA Update
  • Work plan going forwards

*Didnt get much meaningful data from November. Need to integrate with new trigger logic. Jordan will be there at second week of January and David from First week to mock data challenge when this work will happen

*Ed has an idea to get around FPGA clock jitter. He will work with Jordan and David to implement this

2.1 MiniDAQ - DA

2.2 Channel Mappings

2.3 Controls
  • working with Pierrick to finalise integration into his infrastructure. Removing EPICS and using Pierricks variables instead.
  • Hardware for this is to move away from old controllers which are difficult to use and often discontinued. Rest of MICE uses CAEN controllers and the idea is to move to them, min SW changes
  • need spare back from Strathclyde and maybe order a new one

*Temp monitoring work is going on by David and Chris Heidt to allow temp info to be retrieved and stored (from AFE).

  • DA had a discussion with Paul Rubinov about how to speed up data rates. He has some ideas and will should be able to know ho much work is involved in January. He may have the time and be able to be at RAL to do it in the 6months after Jan.
  • DAQ changes suggest a 20% increase in data taking rate

2.4 Hall Probes - MU

We already have a working system but we will improve the SW and make it more user friendly.

3. LED: EO

  • Update on LED tests
  • When we plugged in as much light as expected wasn't seen. This could be due to warm electronics or some miscalculations. It is thought that not enough light is being input. This can be changed externally.
    *Ed is working to understand how much more light we need and this will happen in January with cold cryostats so that it is ready for March time.

4. Tracker lab - AB, GB, CM
Used and clean.
May have lost half our space to Tim Stanley and the RF.

5. Installation

*Installation Manual - GB MU

  • Schedule
    • Current schedule at Integration Schedule - timetable still subject to change (constraints from AFC, etc)
    • Add resource loading? - CM
  • External Waveguides - MU
  • Cryostats - CM, AB, DA
  • Cryostat 2 will be warmed and a new pressure release valve fitted and then they will be pump purge cleaned in January wh they are moved to the pit in
    Cryo 6.07 and cryo 4 6.00

6. Services in the MICE hall: CM

*will need top map which external waveguides can go where.

7. Step IV Analysis Readiness - MU

  • Tracker alignment - MU - ongoing
  • Step4 commissioning in progress.

8. Risks

8.1 Hardware

  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Northside PRY (not us, but knock effect to us)
  • Power supply (spare arrives mid Jan) - MITIGATED

8.2 Software

  • Not yet able to test against multiple tracks per trigger with MC -Action for Durga
  • Real data unpacking broken DA -Ongoing

8.3 Tracker Online Emergency Procedure - MU SB

  • Provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks

A lot of this is almost certainly going to be a work in progress, but it would be good to get it started. At the very least the emergency shutdown and power on instructions and checklist should be available by end of the year. We'll need a lot of this for the Jan 21st mock data run anyway.

  • First thing to do determine list of could fatally fails. - At face to face meeting in January.

9. AoB and summary of actions.

Face to face meeting date January 14th from 2pm at RAL CR9 or 10 TBC.
  • Meeting to discuss: commissioning status and plans, Mock data challenge, possibility of early beam and preparations for it, Step IV readiness, to do list and dates.
  • GB to add all documentation to mice Mine - GB
  • MU to look in what to release as MICE notes and how - MU
  • Kalman doesnt work on straights. - Chris and Adam looking to fix
  • Not yet able to test against multiple tracks per trigger with MC - Durga
  • Real data unpacking broken - DA
  • Face to face agenda -MU
  • Make mock data challenge wiki page for prep discussion. - MU

9.1 AoB

9.2 Revised Actions

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English

Updated by George, Melissa over 8 years ago ยท 2 revisions